OpenStudy (nali):

How would you prepare 175.0mL of 0.150M AgNO3 solution starting with pure solute?

OpenStudy (nali):

@oOKawaiiOo

OpenStudy (nali):

ok can u make sure from my steps first i should find the volume of the solution

OpenStudy (ookawaiioo):

i will show you let me write it down. hold on

OpenStudy (nali):

ok

OpenStudy (ookawaiioo):

\[1000 ml = 1L\] \[175.0ml = 0.175\] \[0.175 L \times \frac{ .150 }{ L } = 0.02625 moles \times \frac{ 170g }{ moles } = 4.4625 grams\] of pure AgNO3